# FloydNet: A Learning Paradigm for Global Relational Reasoning

### v1

@@ B.4 Comparison with Transformers and the Canonical Ordering Problem
Sequence models like the Transformer excel when given a clear, fixed order. Theoretically, any graph problem could be converted into a sequence task if one could define a canonical ordering of its nodes, maintaining a consistent sequence representation for any given graph. However, finding such an ordering is at least as difficult as the Graph Isomorphism (GI) problem, which is not known to be solvable in polynomial time. Graph Transformers attempt to work around this by using positional encodings (PEs), but this is an approximate solution that tries to inject the lost ordering and structural information back into a model that is not inherently designed for it.
@@
B.5 A New Scaling Paradigm: From Large Models to Computationally Intensive Models
@@ Appendix C Formal Discussion on FloydNet’s Expressiveness
The k-FloydNet architecture exactly matches the definition k-Folklore Weisfeiler-Lehman (k-FWL) algorithm, thus it can distinguish non-isomorphic graphs that can be distinguished by the k-FWL test, and vice versa.
@@
The expressive power of a GNN can be characterized by its analogous color refinement scheme. We consider the k-FWL test which is known to have the equivalent power as (k+1)-WL test. The k-FWL test algorithm refines the color of super-edges $\vec{e}=(v_{1},v_{2},\dots,v_{k})$ based on the multiset of k-tuples of colors of neighbor super-edges $\vec{e^{\prime}}=(\vec{e}[i]\leftarrow p)=(v_{1},v_{2},\dots,v_{i-1},p,v_{i+1},\dots,v_{k-1},v_{k})$, which differs to $\vec{e}$ by replacing the $i$-th node to node $p$. We use the notation $\![\cdot$$\cdot$]\!] to denote a tuple, and the notation $\{\!\{\cdot\}\!\}$ to denote a multiset, the formula of the color refinement iteration can be written as:
@@
    The output hash value will change non-linearly with respect to the any of the input values.
@@
k-FloydNet operates on relationship tensor ${\bm{R}}\in\mathbb{R}^{N^{\times k}\times d_{r}}$, which can be view as colors of $N^{k}$ super-edges, and the color is embedded in the continuous space $\mathbb{R}^{d_{r}}$. The Pivotal Attention module will project ${{\bm{R}}}_{\vec{e}}$ to Query tensor $Q_{\vec{e}}$, and project ${{\bm{R}}}_{\vec{e},p,i}={{\bm{R}}}_{\vec{e}[i]\xleftarrow{}p}$ to Key tensor $K_{\vec{e},p,i}$ and Value tensor $V_{\vec{e},p,i}$ for each pivot $p\in V$ and index $i\in[1,k]$. Different linear projection modules are applied for each index $i$ (as shown in Figure˜1). Then a MultiheadAttention is applied to these Query, Key, and Value tensors to get the final output tensor ${{\bm{R}}^{\prime}}_{\vec{e}}$, which we view as the refined color.
@@
    The output embedding ${{\bm{R}}^{\prime}}_{\vec{e}}$ will change non-linearly with respect to the any of the input embeddings ${\bm{R}}_{\vec{e^{\prime}}}$ due to the MLPs, which can approximate any continuous function to an arbitrary degree of accuracy, according to the Universal Approximation Theorem.
So Pivotal Attention with color represented by ${{\bm{R}}}_{\vec{e}}\in\mathbb{R}^{d_{r}}$ is an valid implementation of the color refinement iteration of the k-FWL test algorithm. The k-FloydNet consists of multiple layers of Pivotal Attention, reflecting the multiple iterations of the color refinement process. Thus k-FloydNet is a k-FWL test algorithm.
@@ D.3 Details on the TSP Benchmark
  • Concorde: Concorde(Aldous and Percus, 2003) is widely regarded as the most effective exact TSP solver for general instances. It uses a sophisticated branch-and-cut algorithm built upon integer linear programming. While it can find the provably optimal solution, its runtime can be exponential in the worst case. We use Concorde to generate the ground-truth labels (the optimal tour) for our dataset.

### v2

@@ B.5 Comparison with Transformers and the Canonical Ordering Problem
Sequence models like the Transformer excel when given a clear, fixed order. Theoretically, any graph problem could be converted into a sequence task if one could define a canonical ordering of its nodes, maintaining a consistent sequence representation for any given graph. However, finding such an ordering is at least as difficult as the Graph Isomorphism (GI) problem, which is not known to be solvable in polynomial time (Babai, 2016; Grohe, 2017; McKay and Piperno, 2014). Graph Transformers attempt to work around this by using positional encodings (PEs), but this is an approximate solution that tries to inject the lost ordering and structural information back into a model that is not inherently designed for it.
@@
B.6 A New Scaling Paradigm: From Large Models to Computationally Intensive Models
@@ Appendix C Formal Discussion on FloydNet’s Expressiveness
The k-FloydNet architecture exactly matches the definition of the k-Folklore Weisfeiler–Lehman (k-FWL) algorithm. Therefore, it distinguishes exactly the same pairs of graphs as the k-FWL test (and vice versa).
@@
The expressive power of a GNN can be characterized by its analogous color refinement scheme. We consider the k-FWL test, which is known to be equivalent in distinguishing power to the (k+1)-WL test. In k-FWL, the color of each super-edge $\vec{e}=(v_{1},v_{2},\dots,v_{k})$ is refined using the multiset of k-tuples obtained by replacing one position with a pivot $p\in V$: for each index $i\in[1,k]$, define the neighbor super-edge $\vec{e}[i]\xleftarrow{}p=(v_{1},\dots,v_{i-1},p,v_{i+1},\dots,v_{k})$. We use $\![\cdot$$\cdot$]\!] to denote an ordered tuple and $\{\!\{\cdot\}\!\}$ to denote a multiset; the refinement update can be written as:
@@
    The output hash value will change non-linearly with respect to any of the input values.
@@
k-FloydNet operates on a relationship tensor ${\bm{R}}\in\mathbb{R}^{N^{\times k}\times d_{r}}$, which can be viewed as (continuous) colors of the $N^{k}$ super-edges. The Pivotal Attention module projects ${\bm{R}}_{\vec{e}}$ into a Query $Q_{\vec{e}}$, and projects ${\bm{R}}_{\vec{e}[i]\xleftarrow{}p}$ into a Key $K_{\vec{e},p,i}$ and Value $V_{\vec{e},p,i}$ for each pivot $p\in V$ and index $i\in[1,k]$. Different linear projection modules are applied for different $i$ (as shown in Figure˜1). A MultiheadAttention then aggregates over pivots to produce an output ${\bm{R}}^{\prime}_{\vec{e}}$, which we interpret as the refined color.
@@
    The output embedding ${\bm{R}}^{\prime}_{\vec{e}}$ will change non-linearly with respect to any input embedding ${\bm{R}}_{\vec{e^{\prime}}}$ due to the use of MLPs, which can approximate any continuous function to arbitrary accuracy (Universal Approximation Theorem).
Therefore, Pivotal Attention with colors represented by ${\bm{R}}_{\vec{e}}\in\mathbb{R}^{d_{r}}$ provides a valid implementation of a k-FWL-style refinement step. Stacking multiple Pivotal Attention layers corresponds to iterating this refinement, hence k-FloydNet implements the k-FWL test.
@@ D.3 Details on the TSP Benchmark
  • Concorde: Concorde(Applegate et al., 1998) is widely regarded as the most effective exact TSP solver for general instances. It uses a sophisticated branch-and-cut algorithm built upon integer linear programming. While it can find the provably optimal solution, its runtime can be exponential in the worst case. We use Concorde to generate the ground-truth labels (the optimal tour) for our dataset.